# High-dimensional covariance estimation by pairwise likelihood truncation

### v1

@@ 2.3 Relationship with adaptive thresholding
Adaptive thresholding sets a covariance estimate to zero if the absolute value of the $z$-score $S_{jk}/\text{SE}_{jk}$ is sufficiently small. In our setting, $\text{SE}_{jk}=\{n\widehat{\text{var}}({\mathbf{u}}_{jk})\}^{-1/2}$ is the standard error, where $\widehat{\text{var}}({\mathbf{u}}_{jk})=n^{-1}\sum_{i=1}^{n}||{\mathbf{u}}_{jk}({\mathbf{s}};{\bm{X}}^{(i)})||^{2}_{2}$ is the empirical Fisher information based on the $jk$th pair. While this approach is widely used for its extreme simplicity in large multivariate problems, $\text{SE}_{jk}$ clearly ignores the information already contributed by other covariance estimates $S_{rs}$, $rs\neq jk$. The following analysis of the first-order conditions associated with the objective (13) reveals that the sparse estimator $\hat{\bm{\theta}}$ is a type of adaptive thresholding estimator, with standard errors adjusted sequentially based on the information of the other non-zero covariance estimates.
Inspecting the non-zero elements $\hat{\mathbf{w}}_{\hat{\mathcal{E}}}$ in the solution (15) provides further insights into the selection process. Consider first the case where $\hat{w}_{jk}\neq 0$; for this to be true we must have that
@@
Taking absolute value on both sides of (22) shows that a sufficient condition for $\hat{w}_{jk}\neq 0$ for some $j<k$ is
@@
The adjusted standard error expression in (24) emphasizes an important difference between adaptive thresholding method and our approach. Particularly, note that $\widehat{\text{cov}}({\mathbf{u}}_{jk},\hat{\mathbf{w}}^{\top}{\mathbf{u}})$ is the covariance between the $jk$th score and the pairwise likelihood based on scores different from ${\mathbf{u}}_{jk}$. This means that, differently from the usual standard error formula $\text{SE}_{jk}=\{n\widehat{\text{var}}({\mathbf{u}}_{jk})\}^{-1/2}$, the adjusted standard error $\text{SE}^{\text{adj}}_{jk}$ is computed by removing the portion of information already explained by linear combination of currently selected scores.
@@ 2.4 Selection of $\lambda$
by Slutsky’s Theorem. The alternative hypothesis $H_{1}$ is suggested if $nS^{2}_{jk}/(S_{jk}^{2}+S_{jj}S_{kk})>q_{\alpha}(1)$, where $q_{\alpha}(1)$ refers to the $(1-\alpha)$ quantile of $\chi^{2}(1)$ distribution and $\alpha$ is user-specified. In this case, we should allow nonzero estimate $\hat{\theta}_{jk}$ of $\theta_{jk}$, i.e. $jk\in\hat{\mathcal{E}}$. This leads to the following rule:
@@
where $\gamma=q_{\alpha}(1)$ for some user specified siginificance level $\alpha$.
By the nature of the quadratic optimization problem with $L_{1}$-penalty, in general, more $w_{jk}$ are allowed to be nonzero when $\lambda$ decreases. Criterion (26) can be interpreted as choosing the smallest $\lambda$ to allow as many as possible nonzero elements in $\hat{\bm{\theta}}$ while ensuring all of them are tested significantly different from 0. We distinguish the selection procedure from classical $L_{1}$-penalized optimizations where $\lambda$ is a tuning parameter, by noting that $\lambda$ here is a random variable once $\gamma$ is determined.
@@ 2.5 Cooordinate descent computing
$\hat{w}^{[t+1]}_{jk}=\frac{1}{\hat{\mathbf{J}}_{jk,jk}}\mathcal{S}\left(\hat{\mathbf{J}}_{jk,jk}-\sum_{rs\neq jk}\hat{w}^{[t]}_{rs}\hat{\mathbf{J}}_{jk,rs};\frac{\lambda}{S^{2}_{jk}}\right),$ (27)
while keeping all the other elements $\hat{w}^{(t)}_{rs}$, $rs\neq jk$, fixed. Here $\mathcal{S}(x;\lambda)$ denotes the soft-thresholding operator $\mathcal{S}(x;\lambda)=\text{sign}(x)(|x|-\lambda)_{+}$, and $x_{+}=\text{max}\{0,x\}$, with $\text{sign}(x)$ being the sign function taking values $-1$, 00 and $1$ if $x<0$, $x=0$ and $x>0$ respectively. Operationally, the selection criterion for $\lambda$ described by Equation (26) is implemented through a golden search rule, a univariate numerical optimization method effective in finding a minimum for unimodal functions on a specific interval. Through iterative steps, the procedure progressively narrows down the search interval for $\lambda$ until convergence. Our numerical experiments have shown that this approach is more precise and computationally efficient compared to costly top-down grid searches.
@@ 3 Properties
In this section, we investigate the asymptotic properties of the selection procedure and the behaviour of the implied estimator $\hat{\bm{\theta}}$ of ${\bm{\theta}}$ and of the true covariance matrix ${\bm{\Theta}}$ defined in (12). We simplify notations by letting ${\mathbf{J}}={\mathbf{J}}({\bm{\theta}})$ the covariance of scores at the true value, ${\mathbf{h}}=\text{diag}({\mathbf{J}})$ and $\hat{\mathbf{h}}=\text{diag}(\hat{\mathbf{J}})$. We use $\|A\|_{1}$, $\|A\|_{\infty}$ and $\|B\|_{max}$ to denote the $L_{1}$ norm, infinity norm and the max norm of vecotr $A$ and matrix $B$. Define $\xi=\{jk:\theta_{jk}\neq 0\text{ and }j<k\}\cup\{jj:\theta_{jl}\neq 0\text{ for some }l\neq j\}$ as the index set of the nonzero covariance and their corresponding margins. Let ${\hat{\xi}}=\{jk:\hat{\theta}_{jk}\neq 0\text{ and }j<k\}\cup\{jj:\hat{\theta}_{jl}\neq 0\text{ for some }l\neq j\}$ be the estimate of $\xi$.
Let $m_{0}=|{\mathcal{E}}\setminus\{jj,j=1\ldots,p\}|$ be the number of nonzero bivariate covariances in ${\bm{\theta}}$. We consider two main scenarios: the first, where the data dimension $p$ is fixed, but no particular distributional assumption on ${\bm{X}}$ is made; the second, where $p$ is allowed to grow faster than the sample size $n$ but requires some restrictions on the tail behaviour of the data generating process. In the latter setting, we also allow $m_{0}$ to grow with $p$. For the results presented in this section, we require the following regularity conditions.
@@
Condition C1 prescribes a regular behaviour of the target covariance matrix ${\bm{\Theta}}$ needed when $p$ diverges, while Condition C2 guarantees the existence of ${\mathbf{J}}_{\xi}^{-1}{\mathbf{h}}_{\xi}$, i.e. the minimizer of the ideal population objective $d_{\lambda}({\mathbf{w}};{\bm{\theta}})$, at $\lambda=0$, limited to those entries of ${\mathbf{w}}$ corresponding to zero entries in ${\bm{\Theta}}$. Condition C3 is a main requirement in all our results and states that the columns of ${\mathbf{J}}_{\xi}$ are linearly independent and any $|\xi|-1$ columns of ${\mathbf{J}}_{\xi}$ are linearly independent of ${\mathbf{h}}_{\xi}$, i.e. any linear combination of these columns cannot represent ${\mathbf{h}}_{\xi}$ up to arbitrary precision. For the high-dimensional setting, we also use Condition C4, which assumes an exponential tail behaviour for the entries of the random vector ${\bm{X}}$. This includes variables with tails much heavier than sub-Gaussian tails.
@@
When $p$ and $m_{0}$ are fixed, and $n$ grows selection consistency occurs under rather mild conditions. First note that Theorem 2 does not make restrictive assumptions on the distribution of the original random vector ${\bm{X}}$ or on its moments. The main requirement here is Condition C3, which plays a role analogous to the Irrepresentable Condition (IC) commonly used in the context of Lasso regression and other similar sparsity-inducing penalization methods; e.g., see Zhao and Yu Zhao and Yu (2006). Our version of the IC asserts that the scores corresponding to zero covariances cannot be represented as linear combinations of the scores for the non-zero covariances. Thus, it ensures that our procedure correctly identifies the true set of non-zero covariances and prevents being misled by correlations between scores.
@@
Theorem 3 shows that the probability of a local false positive error – or Type I error – for an individual covariance entry becomes negligible as $n$ increases. Moreover, the TPL selection procedure allows control of the false positive errors locally in large-sample scenarios, irrespective of the data size $p$, as long as the threshold $\gamma$ diverges quickly enough with $n$.
Next, we turn to selection selection consistency when the data dimension $p$ grows exponentially with the sample size $n$. The next theorem is organized into two main parts. Part (i) states that the probability of the overall false positive rate converges to zero. This property implies sparsistency, i.e. the TPL estimator identifies consistently the zero covariance entries as the sample size increases. Part (ii) addresses the power of our selection method in identifying the true nonzero covariances. Specifically, the probability that all true nonzero covariances are correctly identified approaches 1. Together, these results confirm the selection consistency of the estimator, meaning it reliably distinguishes between zero and nonzero covariances.
@@
A few remarks on this result are in order. Our result hinges on Condition C4, which assumes exponential tail behaviour for the entries of the random vector ${\bm{X}}$ with rate $r$. This assumption is milder and more general compared to other common methods often requiring sub-Gaussian random variables (see, e.g., Bickel and Levina Bickel and Levina (2008a) and Lam Lam and Fan (2009)). Consequently, model selection consistency for the TPL selection applies to a broader range of distributions encountered in practical applications, including Gamma-type distributions and those with Gaussian tails.
Different from Theorem 2 where $p$ and $m_{0}$ are fixed, the quantities $r_{1}>0$ and $r_{2}>0$ determine the maximum rates allowed for $p$ and $m_{0}$, respectively, with $r_{1}$ being negatively related to $r_{2}$. Lighter tails for the data distribution means that one can afford a larger data dimension $p$. Moreover, it should be easier to handle a larger proportion of non-zero elements when $p$ grows slowly, and vice versa. To capture this behaviour, instead of fixing $r_{1}$ at its maximum, Theorem 4 allows one to specify the trade-off between sparsity and total data dimension by choosing $r_{1}$ within the range $0<r_{1}<\frac{r}{8+r}$, and allowing $r_{2}$ to increase accordingly. For example, if ${\bm{X}}$ is Gaussian, the tail of each component decays at rate $\exp(-t^{2}/2)$ as $t$ increases, meaning that $r=2$ and $50<r_{1}<1/$ 5. At the maximal rate allowed for $p$, $r_{1}=1/5$ and $\gamma$ should be between $n^{0.6}$ and $n^{0.9}$. On the other hand, at the minimal rate when $r_{1}$ is arbitrarily close to zero, $\gamma$ should be between $n^{0.4}$ and $n^{0.9}$.
@@
We conclude this section by studying the behaviour of the estimator implied by our selection procedure. Denote $\hat{\bm{\Theta}}^{oml}$ the oracle ML estimator of the covariance matrix which assumes the knowledge of no-zero entries of ${\bm{\Theta}}$, with elements $\hat{\bm{\Theta}}^{oml}_{jk}=S_{jk}$ if $\theta_{jk}\neq 0$, and $\hat{\bm{\Theta}}^{oml}_{jk}=0$ if $\theta_{jk}=0$. We compare the TPL estimator to $\hat{\bm{\Theta}}^{oml}$.
@@
Under the assumptions of Theorem 4, we have $\Pr(\hat{\bm{\Theta}}={\bm{\Theta}}^{oml})\to 1$, as $n\to\infty$.
Corollary 1 states that as the sample size grows our estimator converges in probability to the oracle ML estimator, as if the exact locations of the nonzero correlations were known in advance, despite the large size of the original covariance matrix and the initially unknown locations of these nonzero correlations. Furthermore, it can be easily shown that $r_{2}$ is always positive, implying that $m_{0}$ is allowed to grow geometrically with $n$ to achieve the optimal rate in the worst-case scenario. In comparison, Lam and Fan Lam and Fan (2009) demonstrated that, due to the control of bias, $L_{1}$-penalized likelihood approaches are limited to an $O(1)$ number of non-zero covariance elements in the worst-case scenario to obtain the optimal rate, regardless of how fast or slow $p$ grows. Nonetheless, since we allow $p$ to grow exponentially with $n$, $m_{0}$ cannot grow at the rate $O(p)$ stated in Lam and Fan Lam and Fan (2009). This constraint depends on our choice of the score covariance estimator $\hat{\mathbf{J}}$ defined in (14), which would become nearly singular under selection consistency, because $m_{0}$ would be much greater than the sample size $n$.

### v2

@@ 2.3 Relationship with adaptive thresholding
Adaptive thresholding sets a covariance estimate to zero if the absolute value of the $z$-score $S_{jk}/\text{SE}_{jk}$ is sufficiently small. In our setting, $\text{SE}_{jk}=\{n\widehat{\text{var}}({\mathbf{u}}_{jk})\}^{-1/2}$, where $\widehat{\text{var}}({\mathbf{u}}_{jk})=n^{-1}\sum_{i=1}^{n}||{\mathbf{u}}_{jk}({\mathbf{s}};{\bm{X}}^{(i)})||^{2}_{2}$ is the empirical Fisher information based on the the $jk$th pair. While this approach is widely used for its extreme simplicity in large multivariate problems, $\text{SE}_{jk}$ clearly ignores the information already contributed by other covariance estimates $S_{rs}$, $rs\neq jk$. An inspection of the first-order conditions associated with the objective (13) reveals that the sparse estimator $\hat{\bm{\theta}}$ is a type of adaptive thresholding estimator, with standard errors adjusted sequentially based on the information provided by other non-zero covariance estimates.
Inspecting the non-zero elements $\hat{\mathbf{w}}_{\hat{\mathcal{E}}}$ in the solution (15) provides further insights on the selection process. Consider first the case where $\hat{w}_{jk}\neq 0$; for this to be true we must have that
@@
Taking the absolute value on both sides of (22) shows that a sufficient condition for $\hat{w}_{jk}\neq 0$ for some $j<k$ is
@@
with the last expression representing the covariance between the $jk$th score and the pairwise likelihood based on scores different from ${\mathbf{u}}_{jk}$.
The adjusted standard error expression in (24) emphasizes an important difference between adaptive thresholding method and our approach. Unlike the usual standard error formula $\text{SE}_{jk}=\{n\widehat{\text{var}}({\mathbf{u}}_{jk})\}^{-1/2}$, the adjusted standard error $\text{SE}^{\text{adj}}_{jk}$ is computed by removing the portion of variance already explained by the linear combination of currently selected scores.
@@ 2.4 Selection of $\lambda$
by Slutsky’s Theorem. The alternative hypothesis $H_{1}:\theta_{jk}\neq 0$ is rejected if $nS^{2}_{jk}/(S_{jk}^{2}+S_{jj}S_{kk})>q_{\alpha}$, where $q_{\alpha}$ refers to the $(1-\alpha)$-quantile of the chi-square distribution with one degree of freedom and $10<\alpha<$ 1 is user-specified. This leads to the following rule:
@@
where $\gamma=q_{\alpha}$ for some user specified significance level $\alpha$.
By the nature of the constrained quadratic optimization problem defined in (13), more $w_{jk}$ are allowed to be nonzero as $\lambda$ decreases. Criterion (26) can be interpreted as choosing the smallest $\lambda$ to allow as many as possible nonzero elements in $\hat{\bm{\theta}}$, while ensuring that they all test significantly differently from 0. Differently from classical $L_{1}$-penalized estimation where $\lambda$ is a tuning parameter, note that here $\lambda$ is a random variable once $\gamma$ is determined.
@@ 2.5 Cooordinate descent computing
$\hat{w}^{[t+1]}_{jk}=\left\{\begin{array}[]{ccc}\frac{1}{\hat{\mathbf{J}}_{jk,jk}}\left(\hat{\mathbf{J}}_{jk,jk}-\sum_{rs\neq jk}\hat{w}^{[t]}_{rs}\hat{\mathbf{J}}_{jk,rs}\right),&&\text{if }j=k\\ \frac{1}{\hat{\mathbf{J}}_{jk,jk}}\mathcal{S}\left(\hat{\mathbf{J}}_{jk,jk}-\sum_{rs\neq jk}\hat{w}^{[t]}_{rs}\hat{\mathbf{J}}_{jk,rs};\frac{\lambda}{S^{2}_{jk}}\right),&&\text{if }j\neq k\end{array}\right.$ (27)
while keeping all the other elements $\hat{w}^{(t)}_{rs}$, $rs\neq jk$ fixed, until a stopping criterion is met. Here $\mathcal{S}(x;\lambda)$ denotes the soft-thresholding operator $\mathcal{S}(x;\lambda)=\text{sign}(x)(|x|-\lambda)_{+}$, and $x_{+}=\text{max}\{0,x\}$, with $\text{sign}(x)$ being the sign function taking values $-1$, 00 and $1$ if $x<0$, $x=0$ and $x>0$ respectively. Operationally, the selection criterion for $\lambda$ described by Equation (26) is implemented through a golden search rule, a univariate numerical optimization method effective in finding a minimum for unimodal functions on a specific interval. Through iterative steps, the procedure progressively narrows down the search interval for $\lambda$ until convergence. Our numerical experiments have shown that this approach is more precise and computationally efficient compared to costly top-down grid searches.
@@ 3 Properties
In this section, we investigate the asymptotic properties of the selection procedure and the behavior of the implied estimator $\hat{\bm{\theta}}$ defined in (12). We simplify notations by letting ${\mathbf{J}}={\mathbf{J}}({\bm{\theta}})$ be the covariance of scores at the true parameter value, ${\mathbf{h}}=\text{diag}({\mathbf{J}})$ and $\hat{\mathbf{h}}=\text{diag}(\hat{\mathbf{J}})$. We use $\|A\|_{1}$, $\|A\|_{\infty}$ and $\|B\|_{max}$ to denote the $L_{1}$ norm, infinity norm and the max norm of vector $A$ and matrix $B$. Define $\xi=\{jk:\theta_{jk}\neq 0\text{ and }j<k\}\cup\{jj:\theta_{jl}\neq 0\text{ for some }l\neq j\}$ as the index set for the nonzero covariances and corresponding margins. Let ${\hat{\xi}}=\{jk:\hat{\theta}_{jk}\neq 0\text{ and }j<k\}\cup\{jj:\hat{\theta}_{jl}\neq 0\text{ for some }l\neq j\}$ be the estimate of $\xi$.
Let $m_{0}=|{\mathcal{E}}\setminus\{jj,j=1\ldots,p\}|$ be the number of nonzero bivariate covariances in ${\bm{\theta}}$. We consider two main scenarios: the first where the data dimension $p$ is fixed, but no particular distributional assumption on ${\bm{X}}$ is made; the second where $p$ is allowed to grow faster than the sample size $n$, but we require restrictions on the tail behavior of the data generating process. In the latter setting, we also allow $m_{0}$ to grow with $p$. For the results presented in this section, we require the following regularity conditions.
@@
Condition C1 prescribes a regular behavior of the target covariance matrix ${\bm{\Theta}}$ needed when $p$ diverges, while Condition C2 guarantees the existence of ${\mathbf{J}}_{\xi}^{-1}{\mathbf{h}}_{\xi}$, i.e., the minimizer of the ideal population objective $d_{\lambda}({\mathbf{w}};{\bm{\theta}})$, at $\lambda=0$, limited to those entries of ${\mathbf{w}}$ corresponding to zero entries in ${\bm{\Theta}}$. Condition C3 is a main requirement in all our results and states that the columns of ${\mathbf{J}}_{\xi}$ are linearly independent and any $|\xi|-1$ columns of ${\mathbf{J}}_{\xi}$ are linearly independent of ${\mathbf{h}}_{\xi}$, that is any linear combination of these columns cannot represent ${\mathbf{h}}_{\xi}$ up to arbitrary precision. For the high-dimensional setting, we also use Condition C4 which assumes an exponential tail behavior for the entries of the random vector ${\bm{X}}$. This includes variables with tails much heavier than sub-Gaussian tails.
@@
When $p$ and $m_{0}$ are fixed and $n$ grows, selection consistency occurs under rather mild conditions. First note that Theorem 2 does not make restrictive assumptions on the distribution of the original random vector ${\bm{X}}$ or on its moments. The main requirement here is Condition C3, which plays a role analogous to that of the Irrepresentable Condition (IC) commonly used in the context of Lasso regression and other similar sparsity-inducing penalization methods for regression; e.g., see Zhao and Yu Zhao and Yu (2006). Our version of the IC asserts that the scores corresponding to zero covariances cannot be represented as linear combinations of the scores for the non-zero covariances. The condition is needed to ensure that our procedure correctly identifies the true set of non-zero covariances while not being misled by correlations between scores.
@@
Theorem 3 shows that the probability of a local false positive error – or Type I error – for an individual covariance entry becomes negligible as $n$ increases. Moreover, the TPL selection procedure controls the false positive errors locally in large-sample scenarios, irrespective of the data size $p$, as long as the threshold $\gamma$ diverges quickly enough with $n$.
Next, we turn to selection consistency when the data dimension $p$ grows exponentially with the sample size $n$. The next theorem is organized into two main parts. Part (i) states that the probability of the overall false positive rate converges to zero. This property implies the sparsistency property, meaning that the TPL estimator identifies consistently the zero covariance entries as the sample size increases. Part (ii) addresses the power of our selection method in identifying the true nonzero covariances. Specifically, the probability that all true nonzero covariances are correctly identified approaches 1. Together, these results confirm the selection consistency of the estimator.
@@
A few remarks on this result are in order. In the increasing $p$ setting selection consistency hinges on Condition C4, which assumes exponential tail behavior for the entries of the random vector ${\bm{X}}$ with rate $r$. This assumption is milder and more general compared to other common methods often requiring sub-Gaussian random variables (see, e.g., Bickel and Levina Bickel and Levina (2008a) and Lam Lam and Fan (2009)). Model selection consistency for the TPL selection is thus understood to apply to a broader range of distributions encountered in practical applications, including Gamma-type distributions and those with Gaussian tails.
Differently from Theorem 2 where $p$ and $m_{0}$ are fixed, the quantities $r_{1}>0$ and $r_{2}>0$ determine the maximum rates allowed for $p$ and $m_{0}$, respectively, with $r_{1}$ being negatively related to $r_{2}$. Lighter tails for the data distribution imply that one can afford a larger data dimension $p$. Moreover, as expected, it should be easier to handle a larger proportion of non-zero elements when $p$ grows slowly, and vice versa. To capture this behavior, instead of fixing $r_{1}$ at its maximum, Theorem 4 allows one to specify the trade-off between sparsity and total data dimension by choosing $r_{1}$ within the range $0<r_{1}<\frac{r}{8+r}$, and allowing $r_{2}$ to increase accordingly. For example, if ${\bm{X}}$ is Gaussian, the tail of each component decays at rate $\exp(-t^{2}/2)$ as $t$ increases, meaning that $r=2$ and $50<r_{1}<1/$ 5. At the maximal rate allowed for $p$, $r_{1}=1/5$ and $\gamma$ should be between $n^{0.6}$ and $n^{0.9}$. On the other hand, at the minimal rate when $r_{1}$ is arbitrarily close to zero, $\gamma$ should be between $n^{0.4}$ and $n^{0.9}$.
@@
We conclude this section by studying the behavior of the estimator implied by our selection procedure. Denote $\hat{\bm{\Theta}}^{oml}$ the oracle ML estimator of the covariance matrix which assumes the knowledge of no-zero entries of ${\bm{\Theta}}$, with elements $\hat{\bm{\Theta}}^{oml}_{jk}=S_{jk}$ if $\theta_{jk}\neq 0$, and $\hat{\bm{\Theta}}^{oml}_{jk}=0$ if $\theta_{jk}=0$. Next, we compare the TPL estimator to $\hat{\bm{\Theta}}^{oml}$.
@@
Under the assumptions of Theorem 4, we have $\Pr(\hat{\bm{\Theta}}=\hat{\bm{\Theta}}^{oml})\to 1$, as $n\to\infty$.
Corollary 1 states that, as the sample size grows, our estimator converges in probability to the oracle ML estimator, as if the exact locations of the nonzero correlations were known in advance, despite the large size of the original covariance matrix and the initially unknown locations of these nonzero correlations. Furthermore, it can be easily shown that $r_{2}$ is always positive, implying that $m_{0}$ is allowed to grow geometrically with $n$ to achieve the optimal rate in the worst-case scenario. In comparison, Lam and Fan Lam and Fan (2009) demonstrated that, due to the control of bias, $L_{1}$-penalized likelihood approaches are limited to an $O(1)$ number of non-zero covariance elements in the worst-case scenario to obtain the optimal rate, regardless of how fast or slow $p$ grows. Nonetheless, since we allow $p$ to grow exponentially with $n$, $m_{0}$ cannot grow at the rate $O(p)$ stated in Lam and Fan Lam and Fan (2009). This constraint depends on our choice of the score covariance estimator $\hat{\mathbf{J}}$ defined in (14), which would become nearly singular under selection consistency, because $m_{0}$ would be much greater than the sample size $n$.